factorize-
a[sup]4[/sup]+a[sup]2[/sup]+25
1
Expert's answer
2011-05-24T05:07:21-0400
To factorize this expression we have to solve the following equation:
a4+ a2 + 25 = 0
D = 1 - 100 = -99
a2 = (-1 + i√99)/2 and a2 = (-1 - i√99)/2
(a2 + 1/2 - i√99/2)(a2 + 1/2 + i√99/2) for further factorization you have to find the square roots of both a2 s.